OpenStudy (anonymous):

A clerk was asked to change a $10 bill. She returned 9 more dimes than nickels and twenty-one more quarters than dimes. How many coins of each did she return?

OpenStudy (anonymous):

ok so far i have 1,000 = 10 dollars in cents.

OpenStudy (anonymous):

im beng too lazy!

OpenStudy (secret-ninja):

x = nickels x + 9 = dimes (x + 9) + 21 = quarters 0.05x + 0.10(x + 9) + 0.25(x + 30) = 10.00 5x + 10(x + 9) + 25(x + 30) = 1000 5x + 10x + 90 + 25x + 750 = 1000 40x + 840 = 1000 40x = 160 .x = 6 nickels x + 9 = 15 dimes x + 30 = 36 quarters So x = 6 nickels.